OpenStudy (anonymous):

Algebra 2 help, Please?? Multiple choice

OpenStudy (anonymous):

\[\frac{ x^2+4x-5 }{ 5x^2-8x+3 }\times \frac{ 20x-12 }{ x^2-6x-55 }\]

OpenStudy (anonymous):

\[A. \frac{ 4 }{ x+5 }\]

OpenStudy (anonymous):

\[B. \frac{ 4(x-1) }{ x+5 }\]

OpenStudy (anonymous):

\[C. \frac{ 4 }{ x-11 }\]

OpenStudy (anonymous):

\[D. \frac{ 4(x+5) }{ x-11 }\]

OpenStudy (anonymous):

You can simplify that to\[\frac{ (x^2+4x-5)(20x-12) }{ (5x^2-8x+3)(x^2-6x-55) }\]Do you know how to simplify from there?

OpenStudy (anonymous):

Foil??

OpenStudy (anonymous):

Yes, and then you can simplify binomials that are both in the bottom and top of the fraction.